I need to know how to use both elimination and substitution methods on the linear equation 8x-3y=-16
5x+2y=-41

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

elimination
8x-3y=-16
5x+2y=-41
========= same as
16x-6y=-32
15x+6y=-123
==========add these
31x=-155
x= -5
-40-3y=-16
-3y=24
y= -8
============
Substitution
2y=-5x-41
y=(-5/2)x-41/2
substitute into other
8x-3(-5/2)x+123/2
8x+7.5x=-123/2 -32/2, because the second term on the right is -16
15.5x=-155/2
31x=-155
x=-5
and y again is -8.

Answer by ikleyn(52788)